### v1

# Leveraging Large Language Models to Detect Missed Peephole Optimizations

Abstract.
By replacing small, suboptimal instruction sequences within programs with a more efficient equivalent, peephole optimization can not only directly optimize code size and performance, but also potentially enables further transformations in the subsequent optimization pipeline. Although peephole optimization is a critical class of compiler optimizations, discovering new and effective peephole optimizations is challenging as the instruction sets can be extremely complex and diverse. Previous methods either do not scale well or can only capture a limited subset of peephole optimizations.
In this work, we leverage Large Language Models (LLMs) to detect missed peephole optimizations. We propose Lampo, a novel automated framework that synergistically combines the creative but unreliable code optimization ability of LLMs with rigorous correctness verification performed by translation validation tools, integrated in a feedback-driven iterative process.
Through a comprehensive evaluation within LLVM ecosystems, we show that Lampo can successfully detect up to 17 out of 25 previously reported missed optimizations in LLVM on average, and that 22 out of 25 can potentially be found by Lampo with different LLMs. For comparison, the state-of-the-art superoptimizer for LLVM, Souper, identified 15 of them. Moreover, within seven months of development and intermittent experiments, Lampo found 26 missed peephole optimizations, 15 of which have been confirmed and 6 already fixed. These results demonstrate Lampo’s strong potential in continuously detecting missed peephole optimizations.
compiler, peephole optimization, large language model
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Software and its engineering Compilers
1. Introduction
Compiler optimizations are a cornerstone of modern software development, designed to enhance a program’s performance without altering its behavior (Aho et al., 2006). A large portion of optimizations are applied in a series of modular stages called compiler optimization passes, which analyze and transform a program’s intermediate representation (IR). Among these, peephole optimization is a critical technique that performs local transformations on small, suboptimal instruction sequences. It operates by identifying a “window” (or “peephole”) of dependent instructions and replacing them with a potentially more efficient and refined (i.e., equivalent or with less non-determinism when the original sequence has undefined behavior) sequence. This process is essential for tasks like eliminating redundant code, performing constant folding, and simplifying complex operations. Importantly, these local improvements can also enable more significant optimizations by subsequent passes, making peephole optimization a foundational and indispensable technique in compiler design.
Despite its importance, the development of a comprehensive, effective peephole optimizer remains a significant challenge. Unlike other optimizations, such as dead code elimination or loop-invariant code motion, which may be implemented with a single, unified algorithm, peephole optimizations are a collection of pattern-matching and rewriting rules. The complexity and vastness of modern instruction sets cause the number of potential optimization patterns to multiply rapidly, making it exceedingly difficult to develop a complete optimizer. This is why optimizers in widely used compilers like LLVM are continuously evolving, with new patterns being generalized or added (Menendez and Nagarakatte, 2016; ins, nd).
Previous approaches for discovering new peephole optimizations can be classified into three categories:
Manual Inspection: A straightforward way to discover new peephole optimization is manually inspecting the generated IR or assembly code to identify suboptimal instruction patterns. This approach has led to the discovery of a considerable number of optimizations; however, it is labor-intensive, requires deep domain expertise, and has limited scalability.
Differential Testing: Prior studies have proposed using differential testing to automatically detect missed optimization opportunities (Theodoridis et al., 2022; Liu et al., 2023; Barany, 2018; Italiano and Cummins, 2024). This is typically performed in one of the two ways: by identifying differences among the results of different optimizing compilers, or by comparing the compilation results of a pair of semantically equivalent programs. While these approaches have proven effective, they are limited to discovering optimizations that have already been implemented in other compilers or that are not performing as expected in certain cases, rather than discovering entirely new classes of missed optimizations.
Superoptimization: Given a sequence of instructions, a superoptimizer aims to find the optimal instruction sequence through stochastic search or program synthesis (Sasnauskas et al., 2017; Schkufza et al., 2013; Joshi et al., 2002; Bansal and Aiken, 2006; Buchwald, 2015; Bansal and Aiken, 2008). The resulting transformation can then be generalized and implemented as a peephole optimization. However, these approaches are typically computationally expensive, which constrains the code size they can process and limits them to supporting only a subset of the instruction set. For example, Souper, a synthesized superoptimizer for LLVM IR, does not support memory, floating-point, or vector instructions.
Lampo.  This paper explores the feasibility of using Large Language Models (LLMs) to discover new peephole optimizations. While LLMs are highly effective at creative exploration and reasoning, they are prone to generating incorrect or “hallucinated” answers. Conversely, formal verification can guarantee the correctness of an optimization but is often unsuited for the creative task of searching for new ones. We propose Lampo, a novel technique that resolves this tension by combining the creative search capabilities of LLMs with the rigorous guarantees of formal verification. The Lampo framework operates as an automated discovery engine for missed peephole optimizations. Its workflow is built around three key components: an extractor, an LLM-based optimizer, and a verifier. First, the extractor analyzes programs to identify candidate instruction sequences for optimization. These candidates are then passed to the LLM-based optimizer, which suggests a potentially more efficient equivalent. The verifier then rigorously checks the proposed optimization for refinement and performance improvement. If the optimization is verified, it is saved for further investigation. If the verification fails, the verifier’s output is used as targeted feedback to the LLM, guiding the LLM to correct its proposal in a new attempt. This closed-loop process allows us to systematically and automatically find new, verified peephole optimizations.
We thoroughly evaluated Lampo’s ability to find new peephole optimizations. We first curated a benchmark suite of 25 missed peephole optimizations that were recently reported in the LLVM repository. Our results show that Lampo can identify up to 22 of them, while Souper, the state-of-the-art superoptimizer, only detected 15. Furthermore, we tested Lampo’s capability to discover unreported, missed optimizations. Within seven months of development and intermittent testing, Lampo found 26 missed peephole optimizations, with 15 confirmed and 6 already fixed in LLVM. Notably, Souper fails to identify 7 of these confirmed cases. These results clearly demonstrate the strong potential of our approach, which combines the creative search of LLMs with the rigorous correctness of formal verification, to continuously find new optimizations as the reasoning abilities of LLMs advance.
Contributions.  We make the following contributions.
  • We present Lampo, a novel framework that synergistically combines the creative search capabilities of LLMs with the rigorous guarantees of formal verification, providing an automated solution to the long-standing challenge of discovering new peephole optimizations.
  • We introduce a new methodology for compiler optimization discovery that leverages LLMs to overcome the limitations of prior approaches like manual inspection, differential testing, and traditional superoptimization. A key innovation of our method is a closed-loop feedback mechanism, where formal verification results are used to iteratively guide the LLM’s search for correct optimizations.
  • We conducted a comprehensive evaluation of Lampo within the LLVM ecosystem, demonstrating its ability to find unreported missed optimizations. Lampo has successfully identified 26 missed peephole optimizations, with 15 confirmed and 6 already fixed in LLVM. Notably, 7 of these optimizations cannot be detected by Souper, the state-of-the-art superoptimizers for LLVM. These results highlight the advantages of Lampo over prior approaches and establish it as a validated, practical tool for compiler development.
2. Background
We built Lampo on top of LLVM and used Lampo to find missed peephole optimizations within LLVM. We chose LLVM because of its flexible and user-friendly tool-chain. While our implementation targets LLVM, the core concept of Lampo is general and can be applied to other compilers. This section introduces necessary background knowledge on LLVM IR, peephole optimizations, and translation validation.
pub fn clamp(inp: &[i32], out: &mut [u8]) {
for (&i, o) in inp.iter().zip(out.iter_mut()) {
*o = i.clamp(0, 255) as u8;
(a) A clamp function that clamps a sequence of signed 32-bit integers to unsigned 8-bit integers.
define i8 @src(i32 %0) {
%2 = icmp slt i32 %0, 0
%3 = tail call i32 @llvm.umin.i32(i32 %0, i32 255)
%4 = trunc nuw i32 %3 to i8
%5 = select i1 %2, i8 0, i8 %4
ret i8 %5
(b) The suboptimal instruction sequence.
define i8 @tgt(i32 %0) {
%2 = tail call i32 @llvm.smax.i32(i32 %0, i32 0)
%3 = tail call i32 @llvm.umin.i32(i32 %2, i32 255)
%4 = trunc nuw i32 %3 to i8
ret i8 %4
(c) The expected optimal instruction sequence.
1
2define void @clamp(){
3
4vector.body:
5 %i = phi i64 [0,%vector.ph],[%i.next,%vector.body]
6 %0 = getelementptr inbounds nuw i32,ptr %inp,i64 %i
7 %1 = getelementptr inbounds nuw i8,ptr %out,i64 %i
8 %2 = getelementptr inbounds nuw i8,ptr %0,i64 16
9 %wide.load = load <4 x i32>,ptr %0,align 4
10 %wide.load7 = load <4 x i32>,ptr %2,align 4
11 %3 = icmp slt <4 x i32> %wide.load,zeroinitializer
12 %4 = icmp slt <4 x i32> %wide.load7,zeroinitializer
13 %5 = tail call <4 x i32> @llvm.umin.v4i32(
14 <4 x i32> %wide.load, <4 x i32> splat (i32 255) )
15 %6 = tail call <4 x i32> @llvm.umin.v4i32(
16 <4 x i32> %wide.load7, <4 x i32> splat (i32 255) )
17 %7 = trunc nuw <4 x i32> %5 to <4 x i8>
18 %8 = trunc nuw <4 x i32> %6 to <4 x i8>
19 %9 = select <4 x i1> %3,
20 <4 x i8> zeroinitializer, <4 x i8> %7
21 %10 = select <4 x i1> %4,
22 <4 x i8> zeroinitializer, <4 x i8> %8
23 %11 = getelementptr inbounds nuw i8, ptr %1, i64 4
24 store <4 x i8> %9, ptr %1, align 1
25 store <4 x i8> %10, ptr %11, align 1
26 %i.next = add nuw i64 %i, 8
27 %12 = icmp eq i64 %i.next, %n.vec
28 br i1 %12, label %middle.block, label %vector.body
29middle.block:
30
31}
(d) The LLVM IR module of the clamp function in Figure 1(a), with the suboptimal instruction sequence highlighted by the dashed box.
Figure 1. (1(a)) and (1(d)) show a Rust function clamp revealing a missed optimization in LLVM, and its corresponding LLVM IR module. (1(b)) and (1(c)) are a pair of simplified LLVM IR functions illustrate the essence of the optimization.
2.1. LLVM Intermediate Representation
LLVM is an open-source compiler framework for various programming languages and architectures (Lattner, 2002; Lattner and Adve, 2004; llv, ndc). It has a low-level, typed IR in static single-assignment form, designed for expressiveness and extensibility.
LLVM IR programs consist of modules. Each module is a translation unit of the input program, and also is the top level container of all other IR objects, such as functions and global variables (llv, ndb). Figure 1(d) illustrates an LLVM IR module (irrelevant instructions, attributes and declarations are hidden for clarity) that contains a function generated from the original Rust function clamp in Figure 1(a). The IR function is organized into basic blocks, such as vector.body (line 4) shown in the figure. Each basic block comprises a sequence of instructions, has a single entry point (optionally denoted by a label), and ends with a terminator instruction that determines which basic block will be executed next, or whether the function will return. (e.g., br and ret instructions) (llv, ndb).
2.2. Peephole Optimization in LLVM
Peephole optimizations identify and replace inefficient patterns with more efficient ones, by examining small instruction windows (McKeeman, 1965). For instance, the condition ab>a+ba-b>a+b can be optimized to $b<0$. Modern peephole optimizations are applied not only to assembly code, but also at higher compilation levels, including abstract syntax tree and IR (Fischer and LeBlanc Jr, 1991).
The InstCombine pass in LLVM is the primary peephole optimization pass (Mukherjee and Regehr, 2024; Menendez and Nagarakatte, 2016), which applies algebraic simplifications by pattern matching and replacing instruction sequences with more efficient alternatives (Lopes et al., 2015). InstCombine also performs canonicalization, transforming instructions into a consistent form. For example, binary operators with constant operands are rewritten to place the constant on the right-hand side. Such canonicalization reduces the pattern search space and facilitates further optimizations.
However, the development of a comprehensive and effective peephole optimizer is extremely challenging (llv, nda; Davidson and Fraser, 1984). Currently, the LLVM compiler contains over 25,000 lines of C++ code dedicated to peephole optimizations, and this number is still growing with new patterns added or existing patterns being generalized (Menendez and Nagarakatte, 2016; ins, nd).
Missed Optimization.  The Rust program in Figure 1(a) and its corresponding LLVM IR in Figure 1(d) demonstrate a reported missed peephole optimization. The instruction sequence highlighted in the rectangle is suboptimal and can be further optimized. After analysis and simplification, the missed optimization can be illustrated with a pair of minimized LLVM functions as shown in Figures 1(b) and 1(c). In this example, the src function clamps the input using x < 0 ? 0 : umin(x, 255), whereas the tgt function achieves the same effect with umin(smax(x, 0), 255). This transformation reduces the instruction count, resulting in a more efficient implementation.
Figure 2. The overall workflow of Lampo. The verification phase is outlined by the dashed line.
2.3. Superoptimization
Superoptimization aims to transform instruction sequences to their optimal implementation by searching the space of possible instructions (Massalin, 1987). A superoptimizer can work either as an online optimizer that performs optimization during compilation, or as an offline optimization generator that helps identifying missed optimization. As the computational cost of superoptimization grows exponentially with the length of the instruction sequence (i.e., the search space), superoptimization is usually restricted within small code regions and is often used for finding missed peephole optimizations (Bansal and Aiken, 2006; Buchwald, 2015; Sasnauskas et al., 2017).
Souper.  Souper (Sasnauskas et al., 2017) is a superoptimizer for LLVM IR to identify more efficient integer-typed instruction sequences by extracting functions from a module and traversing dataflow edges backward from the return instruction. Unlike exhaustive search, Souper employs counterexample-guided synthesis to accelerate the discovery of optimal replacements. However, its applicability is limited, as it does not support memory accesses, floating-point, or vector instructions.
2.4. Translation Validation
Translation validation is a formal technique for verifying that the target code produced by compilers or optimizers correctly implements the source code (Necula, 2000; Pnueli et al., 1998; Lopes et al., 2021). An automatic translation validation process requires a formal specification of “correct implementation”, which is a refinement relation (Pnueli et al., 1998). For instance, the LLVM IR function produced by a transformation pass, such as the InstCombine pass, should display a subset of the behaviors of the original IR function. The transformation can only eliminate non-determinism by refining undefined behaviors, without introducing new ones. When no undefined behavior exists, the refinement relation becomes an equivalence (Lopes et al., 2021).
Alive2.Alive2 (Lopes et al., 2021) is an automated translation validation tool for LLVM. Given a pair of LLVM functions, Alive2 verifies whether the transformation from the source function to the target function is correct (i.e., whether the source is refined by the target). For example, given the LLVM function pair shown in Figures 1(b) and 1(c), Alive2 can prove that the transformation from src to tgt is correct. Conversely, if a transformation is proven to be incorrect, Alive2 provides a counterexample that refutes the refinement relation between the source and the target. Alive2 is widely used in the LLVM community and is integrated into the LLVM development workflow (llv, nda).
3. Methodology
Figure 2 shows the overall workflow of Lampo. Given a corpus of programs in LLVM IR that have been compiled and optimized by LLVM, Lampo first traverses the basic blocks in these IR programs and extracts all the dependent instruction sequences within each basic block (step 1). For each instruction sequence, Lampo prompts an LLM to output the optimal instruction sequence if the original instruction sequence is considered suboptimal (step 2). However, LLMs are not always reliable. The candidate produced by LLMs can, for example, (1) have syntax error, (2) be sub-optimal, (3) be not canonical, (4) have no refinement relation to the original instruction sequence, or (5) be not interesting as it is identical to the original sequence or is less efficient. To tackle this limitation of LLMs, we incorporate a verification phase as outlined by the dashed line to check all these mistakes that can be made by the LLM (steps 3, 4, and 5). If the instruction sequence proposed by the LLM has syntax error or the optimization is proven to be incorrect, Lampo uses the syntax error message or a counterexample provided by the correctness verifier as feedback, and prompts the LLM to output a correct version based on the feedback until the reattempt limit is reached (steps 6). In our implementation, we set the limit to two. Eventually, if the result generated by the LLMs is interesting (discussed in section 3.3) and the optimization is proven to be correct, Lampo saves the pair of instruction sequences as a potential missed optimization, which is then further analyzed and reported to developers (step 7). The remainder of this section details the workflow.
define <4 x i8> @src(i64 %a0, ptr %a1) {
entry:
%0 = getelementptr inbounds nuw i32,ptr %a1,i64 %a0
%wide.load = load <4 x i32>, ptr %0, align 4
%3 = icmp slt <4 x i32> %wide.load, zeroinitializer
%5 = tail call <4 x i32> @llvm.umin.v4i32(
<4 x i32> %wide.load, <4 x i32> splat (i32 255) )
%7 = trunc nuw <4 x i32> %5 to <4 x i8>
%9 = select <4 x i1> %3,
<4 x i8> zeroinitializer, <4 x i8> %7
ret <4 x i8> %9
(a) An instruction sequence (wrapped as a function) extracted by Lampo from the LLVM IR module shown in Figure 1(d). The dashed box highlights the instructions to be optimized.
define <4 x i8> @src(i64 %a0, ptr %a1) {
entry:
%0 = getelementptr inbounds nuw i32,ptr %a1,i64 %a0
%wide.load = load <4 x i32>, ptr %0, align 4
%smax_0 = smax <4 x i32> %wide.load, zeroinitializer
%smin_255 = smin <4 x i32> %smax_0, splat (i32 255)
%result = trunc nuw <4 x i32> %smin_255 to <4 x i8>
ret <4 x i8> %result
(b) The syntactically invalid candidate initially output by Gemini2.0T.
error: expected instruction opcode
%smax_0 = smax <4 x i32> %wide.load, zeroinitializer
(c) The error message output by opt.
define <4 x i8> @src(i64 %a0, ptr %a1) {
entry:
%0 = getelementptr inbounds nuw i32,ptr %a1,i64 %a0
%wide.load = load <4 x i32>, ptr %0, align 4
%smax_val = tail call <4 x i32> @llvm.smax.v4i32(
<4 x i32> %wide.load, <4 x i32> zeroinitializer
%smin_val = tail call <4 x i32> @llvm.smin.v4i32(
<4 x i32> %smax_val, <4 x i32> splat (i32 255)
%result = trunc nuw <4 x i32> %smin_val to <4 x i8>
ret <4 x i8> %result
(d) The correct candidate produced by Gemini2.0T after receiving the error message as feedback. The dashed box highlights the new instruction that replaces the original icmp and select instructions.
Figure 3. An illustrative example demonstrating the workflow of Lampo.
3.1. Illustrative Example
We use the real-world LLVM module in Figure 1(d) (section 2) to illustrate the workflow of Lampo. Given this module, Lampo first traverses all the basic blocks in each function of this module and extracts instruction sequences from them. Figure 3(a) shows one of the instruction sequences extracted from the basic block vector.body.
Lampo then provides the extracted instruction sequence to an LLM and prompts it to produce a correct and optimal instruction sequence. In this example, Lampo uses gemini-2.0-flash-thinking-exp-01-21 (referred to as Gemini2.0T) as its underlying LLM, and the initial candidate produced by this model, shown in Figure 3(b), contains a syntax error.
After obtaining this candidate, Lampo initiates the verification process. It first invokes the LLVM optimizer opt to check candidate’s syntax and to potentially further optimize and canonicalize the candidate. In this case, since the candidate contains a syntax error, opt outputs an error message as shown in Figure 3(c). Lampo then feeds this error message back to the Gemini2.0T, prompting it to produce a new candidate. Figure 3(d) shows the corrected candidate generated by Gemini2.0T after receiving the feedback.
With this new candidate, Lampo again invokes opt, and now the syntax check passes. Meanwhile, no further optimization and canonicalization is performed by opt since the candidate is already optimal and canonical. The subsequent interestingness checking (discussed in section 3.3) also passes, as the candidate contains one fewer instructions than the original instruction sequence. Finally, Alive2 formally confirms the correctness of the proposed optimization, and Lampo records the original instruction sequence and the candidate as a potential missed optimization.
Souper cannot detect this missed optimization because it does not support the LLVM intrinsic group llvm.umin.*.
3.2. Extracting Instruction Sequences
This subsection introduces the motivation and the design of the extractor (step 1).
Motivation.  The motivation for extracting instruction sequences from an entire LLVM IR module is threefold. First, an LLVM IR module can be extremely large and complex, and from our experience, existing LLMs struggle to identify missed peephole optimizations directly from a complete module. By prompting LLMs to focus on optimizing small instruction sequences, we increase the likelihood of discovering potential missed optimizations. Second, extracting instruction sequences from each IR module allows Lampo to exclude irrelevant information and deduplicate sequences that appear multiple times in the corpus, and thereby increase the overall efficiency and save the cost for LLM inference. Finally, slicing modules into instruction sequences facilitates subsequent verification and analysis by restricting the scope to a small instruction sequence in each iteration.
Input : module, an LLVM module
Input : dedup_set, a set that contains all the seen encoded instruction sequences
Output : A set of unique instruction sequences wrapped to LLVM IR functions
1
2result $\leftarrow\varnothing$
3 foreach function $\in$ module do
4  foreach  basic block bb $\in$ function do
5     $\textnormal{{seq\_set}}\leftarrow\textnormal{{ExtractSeqsFromBB(}}\textnormal{\emph{{{bb}}}}\textnormal{{)}}$
6     foreach seq $\in$ seq_set do
7        wrapped_func $\leftarrow$ WrapAsFunc(seq)
8        if wrapped_func can be further optimized then
9          continue
10        digest $\leftarrow$ Hash(wrapped_func)
11        if digest $\in$ dedup_set then continue
12        $\textnormal{{dedup\_set}}\leftarrow\textnormal{{dedup\_set}}\cup\{\textnormal{{digest}}\}$
13        $\textnormal{{result}}\leftarrow\textnormal{{result}}\cup\{\textnormal{{seq}}\}$
14
15
16
17return result
18
19Function ExtractSeqsFromBB($b$):
20  $\textnormal{{seq\_set}}\leftarrow\varnothing$
21  foreach instruction inst $\in$ bb in reverse order do
22     if inst is a terminator instruction then continue
23     added $\leftarrow$ false
24
25    $\textnormal{{new\_set}}\leftarrow\varnothing$
26     foreach instruction sequence $\textnormal{{seq}}\in\textnormal{{seq\_set}}$ do
27        if any instruction in seq depends on inst then
28           $\textnormal{{new\_set}}\leftarrow\textnormal{{new\_set}}\cup\{[\textnormal{{inst}}]+\textnormal{{seq}}\}$
29           $\textnormal{{added}}\leftarrow\textnormal{{true}}$
30        else  $\textnormal{{new\_set}}\leftarrow\textnormal{{new\_set}}\cup\{\textnormal{{seq}}\}$
31
[…]
33     $\textnormal{{seq\_set}}\leftarrow\textnormal{{new\_set}}$
34
35 return seq_set
36
Algorithm 1 Extracting Instruction Sequences from an LLVM Module – Extract(module, dedup_set)
Algorithm 1 presents the design of Lampo’s extractor, whose goal is to obtain all unique dependent instruction sequences from each basic block in every LLVM IR module in the corpus. The extractor takes two inputs: an LLVM IR module and a set, dedup_set containing all previously seen encoded instruction sequences for deduplication. It outputs a set of unique instruction sequences extracted from the given module. Each of the instruction sequences is wrapped as an LLVM function and passed to the LLM to explore potential optimization opportunity within it. The detailed steps are described below.
Extracting From Basic Blocks.  Given an LLVM IR module, Lampo first initializes result to store the extracted sequences (Algorithm 1). It then traverses each basic block bb of each function in the module (Algorithms 11), and extracts all dependent instruction sequences contained within bb by calling function ExtractSeqsFromBB (Algorithm 1). Specifically, ExtractSeqsFromBB traverses the instructions in bb in reverse order (Algorithm 1). For each instruction inst, if it is a terminator (i.e., the last instruction in a basic block), then it is skipped (Algorithm 1), since terminators only define control flow jumps, and Lampo currently targets only missed optimization within basic blocks. On Algorithm 1, Lampo initializes a flag added as false to indicate whether inst has been inserted into any sequence, and initialize a new set, new_set, which is eventually used to update seq (Algorithm 1). Next, for each instruction sequence seq in seq_set, if any instruction in seq depends on inst, Lampo prepends inst to seq and sets added to true (Algorithms 11). Otherwise, seq remains unchanged (Algorithm 1). If inst is not added to any existing sequences—meaning no traversed instruction depends on instLampo creates a new sequence containing only inst and adds the new sequence to new_set (Algorithms 11).
Verifying and Deduplicating Instruction Sequences.  After Lampo traverses all the instructions within a basic block, a set of instruction sequences can be obtained. For each extracted instruction sequence, Lampo wraps it as an LLVM IR function (Algorithm 1). Specifically, Lampo appends a return instruction that returns the value produced by the last instruction to the sequence and treats all undefined operands in the sequence of instructions as function arguments. With the wrapped function obtained, Lampo checks that LLVM cannot further optimize the instruction sequence (Algorithm 1). This check is necessary because a sequence that is unoptimizable in its original context may become optimizable once isolated. For example, an instruction in the sequence might be used not only by subsequent instructions within the sequence but also by instructions in later basic blocks; once those external uses are removed, the instruction could be combined with others in the sequence. Finally, Lampo computes a hash of the function based on the opcode and operands of each instruction (Algorithm 1), and add instruction sequences whose hashes have not been recorded in dedup_set (Algorithms 11).
3.3. Verifying Outputs of the LLM
After extracting instruction sequences from the corpus and wrapping them as LLVM IR functions (we use the term instruction sequence and function interchangeably in the following), Lampo begins prompting the specified LLM to optimize each function. Each LLM-generated candidate is then verified through three steps.
Preprocessing with opt.  First, Lampo feeds the candidate to opt, the LLVM optimizer, with the -O3 flag to perform aggressive optimization (step 3). The step serves two purposes. First, opt checks whether the function proposed by the LLM is syntactically correct. If the function contains syntax errors, opt outputs an error message, which Lampo can then use as feedback to prompt the LLM to fix the error (step 6). Second, although the LLM may produce a modified version of the original function, the proposed function may not be canonical or optimal w.r.t. opt. In such cases, running opt can further optimize and canonicalize the proposed function reducing the effort required for subsequent manual analysis.
Checking Interestingness.  If the function proposed by the LLM is syntactically correct, Lampo next checks whether the function—after preprocessed by opt — potentially manifests a beneficial optimization. Determining whether the transformation from the original function to the proposed one is beneficial is tricky, because (1) comparison between two functions is not straightforward, as multiple aspects can be considered (e.g., code size and performance on a specific target); (2) even if the proposed function appears significantly better, implementing the corresponding optimization might not be beneficial overall if it hinders other, more valuable optimizations. Due to these challenges, this step is designed to only filter out as many uninteresting cases as possible (i.e., those unlikely to manifest a beneficial optimization), thereby reducing the effort required for subsequent manual analysis. In our prototype, two metrics are considered when checking the interestingness, the instruction count and the total cycles measured by llvm-mca  [^1] A performance analysis tool to statically measures the performance of machine code in a specific CPU based on LLVM’s scheduling models. on a specific target with a specific CPU (e.g., with the target triple set to x86_64-unknown-unknown and the CPU set to btver2) Functions with fewer instructions or fewer total cycles are considered interesting and thus pass the check. Additionally, functions with the same instruction count and total cycles are also considered interesting if they differ syntactically from the original, since such transformations—though not immediately improving the function—may enable further optimization. Interestingness checking is performed before correctness verification, as it is generally more efficient, avoiding the costlier correctness checking for unpromising cases and thus improving overall workflow efficiency.
Verifying Correctness.  If the function proposed by the LLM passes the interestingness checking, Lampo then verifies the correctness of the transformation from the original function to the proposed one using Alive2 (Lopes et al., 2021). If Alive2 proves the transformation incorrect, it provides a counterexample to demonstrate the incorrectness, which Lampo uses as feedback to prompt the LLM to propose a new candidate. If Alive2 encounters a fixable error that prevents the proof—for example, mismatched function signatures—the error message is also used as feedback to guide the LLM in avoiding the issue (step 6). Otherwise, if Alive2 proves the transformation correct, Lampo records the pair of functions as a potential missed optimization (step 7)
4. Evaluation
To comprehensively evaluate Lampo, we conducted experiments to investigate the following research questions.
RQ1: Can Lampo detect previously reported missed optimizations?
RQ2: Can Lampo discover new missed optimizations?
RQ3: How much throughput can Lampo achieve, and what is the associated cost?
4.1. Experimental Setup
Selected LLM Models.Table 1 lists all the models selected for evaluating Lampo. We selected diverse models to investigate how the capabilities of different LLMs affects the effectiveness of Lampo. For RQ1, we selected two open-source and four proprietary models. Specifically, we chose two open-source models with significantly different sizes: Gemma3 (27 billion parameters) and Llama3.3 (70 billion parameters). For proprietary models, we included two base models, i.e., Gemini2.0 and GPT-4.1, and two reasoning models, i.e., Gemini2.0T and o4-mini. These models have relatively early knowledge cut-off dates, allowing us to collect a reasonable number of benchmarks while avoiding potential data leakage. In RQ2, during the long-term, intermittent experiment, we tried different open-source models, and a locally deployed version of llama3.3:70b is the most used one. In RQ3, Llama3.3 and Gemini2.5 (a cost-efficient and low-latency model recommended for scenarios that requires high throughput) are used to evaluate the throughput and cost, covering two common usage scenarios, i.e., using a locally deployed LLM and invoking an LLM via API. For Gemini2.0T and Gemini2.5, we set the reasoning budget to low (i.e., at most 1,024 reasoning tokens).
Table 1. The selected LLMs in evaluation. Gemini2.5 is excluded from RQ1 to prevent potential data leakage.
Model Name Model Version Reasoning Cut-off Date
Gemma3 gemma3:27b No 08/2024
Llama3.3 llama3.3:70b No 12/2023
Gemini2.0 gemini-2.0-flash No 08/2024
Gemini2.0T
gemini-2.0-flash-
thinking-exp-01-21
Yes 08/2024
GPT-4.1 gpt-4.1-2025-04-14 No 06/2024
o4-mini o4-mini-2025-04-16 Yes 06/2024
Gemini2.5 gemini-2.5-flash-lite Yes 01/2025
Baselines.  We used two baselines in our evaluations.
[…]
  • Lampo-. To evaluate the effectiveness of the iterative prompting strategy, we compared Lampo with a variant named Lampo-, which does not prompt the LLM to generate new candidates using feedback (i.e., it omits steps 4 and 7 in Figure 2)
System Configuration.  The experiments were run on an Ubuntu 22.04 server with three NVIDIA RTX 6000 Ada GPUs, Intel Xeon Gold 5217 CPU@3.00GHz, and 384 GB RAM.
4.2. RQ1: Detecting Previously Reported Missed Optimizations
[…]
Issue ID Gemma3 Llama3.3 Gemini2.0 Gemini2.0T GPT-4.1 o4-mini Souper
Lampo- Lampo Lampo- Lampo Lampo- Lampo Lampo- Lampo Lampo- Lampo Lampo- Lampo Default Enum
104875 1 5
107228 5 5 1 4 5
108451 5 5 5 5 4 5 1 4 2 5
108559 5 5 5 5 4 5 1 4 4 5
110591 5 5 5 5 2 5 2 5 3 5
115466 1 1 5 5 5 5 3 4 5 5
118155 3 3 2 2
122235 5 2 5 2 5
122388 5 5 4 4 2 3
126056 1 5 5 1 4 5 5
128475 2 2 4 5
128778 1 5 1 3 5
129947 1
131444
131824 1 3 1 3
132508 2 1 5 5 5 2 2 3 3 5
134318
135411 5 5 1 1 5 5
137161 2 1
141479 5 5 4 5
141753 1 1 2
141930 3 3 4 5 2 2 5 5 5 5
142497 1 1 1
142593 4 2
143259
Average 0.8 1.2 5.2 7 5.8 7.4 10.4 15.6 2.2 7.4 10.0 14.2 N/A N/A
Total 2 3 6 7 7 11 14 21 7 12 14 18 3 14
To quantitatively evaluate the effectiveness of Lampo in finding missed optimization opportunities and to objectively compare it with the baselines, we conducted a controlled experiment to assess how effectively each tool identifies previously reported missed optimizations. For the benchmark suite, we collected 25 missed peephole optimizations from issues in the official LLVM GitHub repository tagged with both missed-optimization and llvm:instcombine labels. To avoid potential data leakage (Zhou et al., 2025; López et al., 2025), we only included issues created after the end of August 2024, which is the latest knowledge cut-off date among all LLMs listed in Table 1 except for Gemini2.5, the model excluded from this research question. For each benchmark, we collected the example suboptimal LLVM IR function provided in the issue and examine whether Lampo, Lampo- and Souper could detect the missed optimization within the function. To mitigate the effects of the inherent non-determinism of LLMs, we repeated each experiment five times for both Lampo and Lampo-. The results of the experiment are shown in Table 2.
Effectiveness of Lampo with Different LLMs.  The effectiveness of Lampo heavily depends on the code understanding and reasoning capabilities of the employed LLM, and LLMs with different capabilities can therefore lead to varying levels of effectiveness. As shown in Table 2, the performance of Lampo varies significantly across different LLMs. With Gemma3, the smallest and weakest of the selected models, Lampo can detect at most 3 out of 25 missed optimizations, and on average only 1.2 per round. With other non-reasoning but more capable models, the performance of Lampo improves substantially: using Llama3.3, Gemini2.0, and GPT-4.1, Lampo can identify up to 7, 11 and 12 missed optimizations, and on average 7.0, 7.4, and 7.4, respectively. Employing reasoning models further enhance the performance of Lampo. With Gemini2.0T and o4-mini, Lampo can detect up to 21 and 18 missed optimizations, and on average 15.6 and 14.2, respectively.
[…]
Lampo vs. Lampo-.  The iterative prompting strategy improves Lampo’s effectiveness in identifying missed optimizations regardless of the model employed. Without iterative prompting, Lampo- failed to identify 1 to 7 optimization that Lampo can detect across different LLMs. Moreover, the average number of optimizations detected by Lampo- is consistently lower than that of Lampo, with the gaps ranging from 0.4 to 5.2.
Table 3. All missed optimizations found by Lampo and reported to LLVM; the Can Souper detect? column indicates whether Souper can detect each of them. The result “timeout” means Souper cannot detect the missed optimization within a 20-minute timeout with Enum set to any value from 1 to 3.
Issue ID Status Can Souper detect?
Default Enum
1 Fixed
2 Confirmed timeout
3 Wontfix
4 Wontfix timeout
5 Fixed
6 Confirmed
7 Confirmed
8 Fixed
9 Fixed
10 Unconfirmed
11 Fixed
12 Unconfirmed
13 Fixed
14 Unconfirmed
15 Confirmed timeout
16 Confirmed
17 Confirmed
18 Unconfirmed
19 Confirmed timeout
20 Confirmed
21 Duplicate
22 Unconfirmed
23 Unconfirmed
24 Unconfirmed
25 Confirmed
26 Unconfirmed
4.3. RQ2: Discovering New Missed Optimization
To demonstrate the practical value of Lampo, we intermittently ran an experiment that employs Lampo to search for new missed optimization in LLVM. In this experiment, the corpus used by Lampo is from an LLVM IR dataset named LLVM Opt Benchmark (Zheng, 2023). This dataset contains optimized IR files from 239 real world projects written in C, C++, or Rust. Since the entire benchmark is too massive to serve as our corpus, we built our corpus by selecting five popular projects for each programming language, i.e., cpython, ffmpeg, linux, openssl, redis, node, protobuf, opencv, z3, pingora, ripgrep, typst, uv, zed. From this corpus, Lampo extracted over 800,000 unique instruction sequences, eliminating around 8.7 million duplicates.
Table 3 lists all missed optimizations found by Lampo and reported to LLVM, and Table 4 presents a summary of reported missed optimizations. Within seven months of development and intermittent testing, Lampo has found 26 potential missed peephole optimizations, with 15 confirmed and 6 already fixed in LLVM. It also detected a bug in Alive2.  [^2] https://anonymous.for.review Among all the reported missed optimizations, we marked 2 of them as wontfix (i.e., Issue 3 and Issue 4 ) according to the feedback from the maintainers. The reasons are: (1) the optimization is typically handled by the backend and (2) the optimization would prevent other valuable optimizations.
Table 4. Summary of Reported Missed Optimizations
Status Reported Confirm Fixed Duplicate Wontfix
Count 26 15 6 1 2
[…]
Case Study 1 (Figures 4(a) and 4(d)).  The src function loads two consecutive 16-bit values from a given address %0 and combines them into a single 32-bit value using zero extension (zext), left shift (shl), and an or instruction. This function can be optimized to directly loading a 32-bit value from the given address. LLVM originally missed this optimization because it did not support merging consecutive loads with different sizes. Specifically, the original unoptimized function loads three consecutive values—two 8-bit values and one 16-bit value. LLVM successfully merges the two 8-bit loads but fails to further merge them with the subsequent 16-bit load. Souper cannot identify this case, as it does not support load and getelementptr instructions.
Case Study 2 (Figures 4(b) and 4(e)).  In the src function, the input %0 is first clamped to be at least 1 by llvm.umax.i8, then doubled with a 1-bit left shift, and finally clamped again to be at least 16. In the tgt function, the initial clamp is removed. This simplification is valid because the second clamp to 16 subsumes the effect of the first clamp, which makes it redundant. Souper cannot identify this missed optimization because it does not support the intrinsic group llvm.umax.*.
Case Study 3 (Figures 4(c) and 4(f)).  In the src function, the input %0 is first checked for being an ordinary (non-NaN) value; if it is NaN, it is replaced with 0.0, otherwise, %0 is kept. The result is then compared against 1.0. In the tgt function, the check and replacement are removed, and %0 is directly compared with 1.0. This simplification is valid because the instruction fcmp oeq (ordered equal) already returns false if %0 is NaN, so explicitly substituting NaNs with 0.0 is redundant. Souper cannot identify this missed optimization as it does not support floating point instructions.
define i32 @src(ptr %0) {
%2 = load i16, ptr %0, align 2
%3 = getelementptr i8, ptr %0, i64 2
%4 = load i16, ptr %3, align 1
%5 = zext i16 %4 to i32
%6 = shl nuw i32 %5, 16
%7 = zext i16 %2 to i32
%8 = or disjoint i32 %6, %7
ret i32 %8 }
(a) Case 1: sub-optimal version.
define i8 @src(i8 %0) {
%2 = call i8 @llvm.umax.i8(
i8 %0, i8 1)
%3 = shl nuw i8 %2, 1
%4 = call i8 @llvm.umax.i8(
i8 %3, i8 16)
ret i8 %4
(b) Case 2: sub-optimal version.
define i1 @src(double %0) {
%2 = fcmp ord double %0, 0.000000e+$0$
%3 = select i1 %2,
double %0, double 0.000000e+$0$
%4 = fcmp oeq double %3, 1.000000e+$0$
ret i1 %4
(c) Case 3: sub-optimal version.
define i32 @tgt(ptr %0) {
%2 = load i32, ptr %0, align 2
ret i32 %2
(d) Case 1: optimized version.
define i8 @tgt(i8 %0) {
%2 = shl nuw i8 %0, 1
%3 = call i8 @llvm.umax.i8(
i8 %2, i8 16)
ret i8 %3 }
(e) Case 2: optimized version.
define i1 @tgt(double %0) {
%2 = fcmp oeq double %0, 1.000000e+$0$
ret i1 %2
(f) Case 3: optimized version.
Figure 4. Three examples of confirmed missed optimizations found by Lampo that Souper is unable to detect.
4.4. RQ3: Throughput and Cost of Lampo
[…]
Tool Lampo Souper
Llama3.3 Gemini2.5 Default Enum =1 Enum =2 Enum =3
Time/Case (s) 26.2 6.7 2.8 37.2 144.4 183.7
# of Timeouts 0 0 0 80 412 616
[…]
For Souper, the default configuration achieved the highest throughput, processing all cases in about 3.9 hours, averaging 2.8 seconds per case. However, as the Enum parameter increases, the throughput decreases sharply. With Enum set to 1, 2, and 3, the execution time rose to 52 hours, 200 hours, and 255 hours, averaging 37.2, 144.4, and 183.7, respectively. Moreover, Souper failed to terminate within 20 minutes in 80, 412, and 616 cases for Enum value of 1, 2, and 3, respectively.
In summary, these results further demonstrate the feasibility and practical value of Lampo. Although Lampo requires additional computation resources for LLM inference compared to Souper, cost-efficient commercial LLMs (e.g., Gemini2.5 from Google) enables Lampo to achieve high throughput with manageable monetary cost. Moreover, as LLMs continue to become faster and more cost-effective, the overall cost-effectiveness of Lampo will likewise improve.
4.5. Threats to Validity
The first threat to validity is potential data leakage. When evaluating the effectiveness of Lampo in finding previously reported missed optimization, it is possible that the training set of the deployed model already contains information related to the missed optimizations, which could inflate the results. To mitigate this threat, we selected LLMs with clear knowledge cut-off dates and only included missed optimizations reported after those dates in our benchmarks. Another concern is that the collected previously reported missed optimizations may not be fully representative. To mitigate this, we gathered as many benchmarks as possible and manually inspected each case to ensure validity. The non-determinism of LLMs poses a threat to the internal validity. We mitigate this by evaluating Lampo with multiple LLMs and repeating the controlled experiment in RQ1 for five times.
5. Related Work
In this section, we discuss the studies related to this work.
Superoptimization.  A superoptimizer takes a sequence of instructions as input and searches for a more efficient, semantically equivalent or refined program. Massalin (Massalin, 1987) first introduced the concept of superoptimization and developed the earliest superoptimizer based on exhaustive enumeration, which, while pioneering, was both inefficient and unsound. To address scalability limitations, subsequent work proposed alternative search strategies, often leveraging SMT solvers for equivalence verification. Joshi et al. developed Denali (Joshi et al., 2002), a goal-directed superoptimizer that employs expert knowledge to encode equality-preserving transformations, thereby constraining the search space. Bansal and Aiken (Bansal and Aiken, 2006) presented a superoptimizer that enumerates only canonical instruction sequences to reduce redundancy and stores computed optimizations in a database to avoid repeated work. Schkufza et al. introduced STOKE (Schkufza et al., 2013), a superoptimizer for x86-64 binaries that formulates superoptimization as a stochastic search problem and applies Markov Chain Monte Carlo (MCMC) sampling to efficiently explore the vast program space. This approach sacrifices completeness but significantly increases the diversity of candidate programs, thereby improving the quality of the resulting code. Sasnauskas et al. proposed Souper (Sasnauskas et al., 2017), a superoptimizer for LLVM IR that extracts integer-typed functions from a module, traverses backward along dataflow edges from the return instruction, and synthesizes more performant equivalents. Souper is designed to be sound, but its applicability is limited to a purely functional, control-flow-free subset of LLVM IR. Hydra (Mukherjee and Regehr, 2024) addresses the limitations of ungeneralized peephole optimizations produced by superoptimizers by generalizing them using program synthesis. In general, superoptimizers are capable of discovering new optimization patterns, but they are computationally expensive, thus often limited to tiny window size or specific instruction sets. In contrast, Lampo leverages LLMs to explore optimization opportunities, which is less restricted by the code size and instruction sets, and thus can potentially detect more diverse missed optimizations.
Differential Testing for Missed Optimizations.  Differential testing (McKeeman, 1998) compares the outputs of two or more programs on identical inputs to detect discrepancies, which may reveal bugs. Barany (Barany, 2018) applied this technique to compilers, identifying missed optimizations by comparing code generated by different compilers for the same source program. Theodoridis et al. (Theodoridis et al., 2022) extended this approach to dead code elimination, while Liu et al. (Liu et al., 2023) used differential testing to evaluate WebAssembly optimizers by comparing their outputs with those of modern C compilers. Italiano and Cummins (Italiano and Cummins, 2024) further advanced differential testing for code size optimizations, proposing methods such as comparing outputs from seed and mutated code, evaluating different optimization settings, and analyzing outputs from different compiler versions or distinct compilers. Differential testing is effective for discovering missed optimizations, but it is limited to optimizations that have been implemented in other compilers or that are not generalized enough. In contrast, Lampo can potentially discover new optimizations that have not been implemented in any compiler.
LLM-Based Approaches for Optimization.  Recent advances in LLMs have demonstrated their promise in program synthesis, including code generation and optimization tasks. A number of studies have explored the use of these models for optimizations (Garg et al., 2023; Gao et al., 2024; Cummins et al., 2023; Grubisic et al., 2024; Italiano and Cummins, 2024). Some of them focus on source-level code optimization and do not address the problem of finding missed optimizations (Garg et al., 2023; Gao et al., 2024). Cummins et al. (Cummins et al., 2023) investigated the use of LLMs to generate optimized LLVM IR and recommend compiler passes, highlighting their potential but also their limitations in terms of reusability and computational cost. Grubisic et al. (Grubisic et al., 2024) incorporated feedback from optimizer-generated code to enhance optimization quality. The only prior work that directly targets missed optimizations with LLMs is by Italiano and Cummins (Italiano and Cummins, 2024). Their approach employs LLMs to mutate code for differential testing, achieving effective results but restricting the discovery of novel optimization patterns. In contrast, Lampo leverages the code understanding and reasoning capabilities of LLMs to directly detect suboptimal code snippets, which can potentially discovery new unimplemented optimization patterns.
6. Conclusion
This paper introduced Lampo, a novel framework that bridges the creative capabilities of LLM with the rigorous guarantees of formal verification to tackle the difficult challenge of discovering new peephole optimizations. Our methodology overcomes the limitations of previous approaches by integrating an LLM-based optimizer within a closed-loop system, where verification feedback iteratively refines the LLM’s search for correct and efficient code transformations. Through a comprehensive evaluation on the LLVM compiler, we demonstrated Lampo’s effectiveness in finding a substantial number of previously unreported optimizations. The fact that many of these have been confirmed and fixed in the official LLVM codebase highlights our framework’s practical significance and its ability to outperform state-of-the-art superoptimizers.
Looking ahead, this research opens up exciting new avenues for the use of LLMs in compiler design and beyond. The success of Lampo suggests that similar hybrid methodologies, which combine the generative power of LLMs with rigorous formal methods, could be applied to other complex, pattern-based optimization problems. As the reasoning abilities of LLMs continue to advance, we believe that frameworks like Lampo will become an indispensable tool for continuously improving the performance of modern compilers, paving the way for more efficient and robust software.
Acknowledgements.
This work was partially funded by an unrestricted gift from Google.